OpenStudy (anonymous):

Help Please dx Write the equation in function notation. y= -2/3x + 400

OpenStudy (campbell_st):

well replace y with f(x)

OpenStudy (anonymous):

i dont understand

OpenStudy (campbell_st):

ok.... here is an equation y = x written using function notation its f(x) = x

OpenStudy (the_fizicx99):

f(x) is another way of saying y, it's a function of x. Being y is dependent on x

OpenStudy (anonymous):

im so lost

OpenStudy (the_fizicx99):

y = f(x) just replace y with f(x). That's it :P

OpenStudy (campbell_st):

here is a link that may explain it http://mrsgalgebra.pbworks.com/f/Alg+1+SOL+A.7+4.7+Function+Notation+Notes+PDF.pdf

OpenStudy (anonymous):

so its f(x) = -2/3 +400 ?

OpenStudy (campbell_st):

well you are missing the independent variable... x

OpenStudy (anonymous):

omfg -.-

OpenStudy (the_fizicx99):

-2/3 has an x :P f(x) = -2/3x + 400 * without the x you can't really do anything

OpenStudy (anonymous):

so thats the answer?

OpenStudy (the_fizicx99):

Yeah o.0 you're just replacing y with the function notation -- f(x)

OpenStudy (anonymous):

ohh

OpenStudy (the_fizicx99):

Yeah :>